SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (17th April, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Brandon Drummond, Al Eisner, Fang Fang, Kim Hojeong, Arik Kreisel, Su Dong.

Operations

  • DCZ crate damage Sunday swing shift from DCZ synchronisation request. ODF log files showed many messages of spurious interrupt, reframe errors and out of synch errors. Pilot power cycled TSF crates.
  • Luminosity reached 10^34 over the weekend - a number of trigger parameters need updating.
    • L1 deadtime alarm threshold is 2% - should be updated to 3% for now but it would be nice to separate out the deadtime due to busy and full, and have an alarm that follows deadtime due to busy.
    • Paged a number of times for increasing numbers of GLT transmission errors and GLT corruption. Should remake JAS plots to display errors as % of events rather than an absolute number.
  • Once beam conditions are more stable, it would be good to take a L1 passthrough run with both on- and off-peak data.
  • Tomorrow is the monthly TRG talk in the 3.45 meeting. Please look through the slides.
  • Debbie will be away from 24th April - 3rd May. Hojeong has kindly agreed to take the pager 26th-3rd, but someone is needed to cover the first couple of days. Any takers?

EMT

  • The JAS plot of M phi distribution shows a peak with high backgrounds (essentially, all the time now). Should make instructions clearer for shifters to say which beam is causing the backgrounds.

DCT/DCZ

  • DCH-DCT calibration is still not working at the teststand. A number of hardaware and software changes have been made on the DCH side but still the 0x4000 error remains and there are no TSF hits.
    • After the meeting we swapped the two TSF-X boards (and swapped them back) and ran a couple of successful calibrations - it seems the problem was with bad connection to the backplate.

Online

  • TSF memory was tested as part of the debug of teststand problems - no problmes were found with each board but it was not possible to test all boards simultaneously. It is believed unrealted to the calibration problems.

Data Quality and Monitoring

  • Run 3 stripcharts can be found here.

Trickle injection monitoring

  • Matt has tested the trickle quality monitoring module and it will be put into L3.

Trigger Configuration

  • We are awaiting the SP9 MC (with 1Zk simulated) to be available to re-do the trigger analyses.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.